Order: acetaminophen elixir 10mg/kg po. Child weight is 10bs. Available: Acetaminophen elixir 160mg/ 5ml. How much Acetaminophen will you administer in ml?order: acetaminophen elixir 10mg/kg po. Child weight is 10bs. Available: Acetaminophen elixir 160mg/ 5ml. How much Acetaminophen will you administer in ml?
Answer #1

Ordered dose= 10mg/kg

Weight of the child =10 lbs . (1 lb= 0.453kg)

So, weight in kg= 10× 0.453=4.53kg

Available dose= 160 mg in 5 ml

Therefore, 1 ml= 160÷ 5=32 mg

So, ordered dose for 4.53 kg child = 10 mg×4.53 kg =45.3 mg

1 ml contains 32 mg

Thus, 45.3 mg = x ml

(Proportions method)

32/1=45.3/x

  • X= 45.3/ 32= 1.415 ml= 1.46 ml

​​​​​​​so for one dose of acetaminophen for 4.53 kg child= 1.46 ml